How to solve the following system of equations simultaneously
y=2x+3
y=7-2x

Answer by Alan3354(69443)

How to solve the following system of equations simultaneously
y=2x+3
y=7-2x
----------
There's more than one method.
For these, since they both equal y, they equal each other.
y=2x+3
y=7-2x
------
2x+3 = 7-2x